OpenStudy (anonymous):

Twelve is the geometric mean of 9 and what number?

OpenStudy (amistre64):

12 = sqrt(9n)

Directrix (directrix):

If 12 is the geometric mean between 9 and a mystery number x, then the following is true: 9 is to 12 as 12 is to x 9/12 = 12/x Cross Multiply and solve for x. @mathloser101